OpenStudy (anonymous):

If a certain nonzero n number is squared, the result equals p times n. The value of n in terms of p is: p. Why?

OpenStudy (anonymous):

certainnonzero^2 = p*n n=p therefore p*p=p^2 ? idk lol

OpenStudy (anonymous):